# Robust Online Conformal Prediction under Uniform Label Noise

### v2

@@ 4.1 Theoretical results
Table 1: Average performance of different methods under uniform noisy labels across 4 score functions, using ResNet18. The performance on each score function is provided in Appendix B.6. “Baseline” denotes the standard online conformal prediction methods. We include two learning rate schedules: constant learning rate $\eta=0.05$ and dynamic learning rates $\eta_{t}=1/t^{1/2+\varepsilon}$ where $\varepsilon=0.1$. $\downarrow$” indicates smaller values are better and Bold numbers are superior results.
@@
Constant $\epsilon=0.05$ Baseline 3.942 2.867 11.93 30.66 4.163 3.289 101.0 231.7
@@
Dynamic $\epsilon=0.05$ Baseline 4.106 2.780 6.527 21.97 4.498 2.584 31.15 128.1
@@
Interpretation.
The proof is detailed in Appendix D.1. Similar to the findings in Proposition 3.1, this result demonstrates that label noise induces a coverage gap given by: $\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\left|\mathcal{C}_{t}(X_{t})\right|\right]\right)$, which quantifies the discrepancy between the actual mis-coverage rate and the target rate $\alpha$. The primary distinction from Proposition 3.1 lies in the constant $B$.
Remark 4.2.
Proposition 4.1 generalizes our previous analysis: the result of constant learning rate in Proposition 3.1 emerges as a special case when $\eta_{t}\equiv\eta$ for all $t\in\mathbb{N}^{+}$. More broadly, this result holds for any sequence of learning rates that satisfies the condition $\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|/T\to 0$ as $T\to+\infty$. Additionally, empirical evidence corroborating our theory can be found in Appendix B.5.
4.2 Method
Our analysis establishes that label noise significantly impacts coverage rate regardless of the learning rate schedule. Building on these insights, we extend our method - NR-OCP to the case of dynamic learning rates. Specifically, we update the threshold by:
$=\hat{\tau}_{t}-\eta_{t}\cdot\nabla_{\hat{\tau}_{t}}l_{1}(\hat{\tau}_{t},\tilde{S}_{t})+\eta_{t}\cdot\nabla_{\hat{\tau}_{t}}l_{2}(\hat{\tau}_{t},\tilde{S}_{t}),$
where
$\nabla_{\hat{\tau}_{t}}l_{1}(\hat{\tau}_{t},\tilde{S}_{t})=\frac{1}{1-\epsilon}\left[\mathds{1}\left\{\tilde{S}_{t}\leq\hat{\tau}_{t}\right\}-(1-\alpha)\right];$
$\nabla_{\hat{\tau}_{t}}l_{2}(\hat{\tau}_{t},S_{t,y})=\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\left[\mathds{1}\left\{S_{t,y}\leq\hat{\tau}_{t}\right\}-(1-\alpha)\right].$
For the convergence of coverage rate, our theoretical results show that, under dynamic learning rate, NR-OCP achieves convergence rates of $\mathrm{EmCovErr(T)}=\mathcal{O}(T^{-1/2})$ and $\mathrm{ExCovErr(T)}=\mathcal{O}(T^{-1/2})$. The proofs are presented in Appendix D.2 and D.3.
Proposition 4.3.
$\mathrm{ExErr(T)}\leq\sqrt{\frac{\log(2/\delta)}{1-\epsilon}}\cdot\frac{1}{\sqrt{T}}+\frac{C}{T}$
where
$C=\left[\left(1+\max_{1\leq t\leq T-1}\eta_{t}\cdot\frac{1+\epsilon}{1-\epsilon}\right)\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|\right].$
Proposition 4.4.
$\mathrm{EmErr(T)}\leq\frac{2-\epsilon}{1-\epsilon}\sqrt{2\log\left(\frac{4}{\delta}\right)}\cdot\frac{1}{\sqrt{T}}+\frac{\gamma}{T}$
where
$\gamma=\left[\left(1+\max_{1\leq t\leq T-1}\eta_{t}\cdot\frac{1+\epsilon}{1-\epsilon}\right)\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|\right]$
with at least $1-\delta$ probability.
In Proposition 4.3 and 4.4, we establish that for any sequence of learning rates that satisfies the condition $\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|/T\to 0$ as $T\to+\infty$, both expected and empirical coverage errors asymptotically vanish with a convergence rate of $\mathcal{O}(T^{-1/2})$. Therefore, by applying our method, the long-term coverage rate would approach the desired target level of $1-\alpha$. In addition, we formally establish the convergence of $\hat{\tau}_{t}$ towards the global minima in Appendix B.2 by extending the standard convergence analysis of stochastic gradient descent.
(a) Constant learning rate
(b) Dynamic learning rate
Figure 2: Performance comparison of various methods under uniform noisy labels with noise rate $\epsilon=0.05$, employing (a) constant learning rate $\eta=0.05$ and (b) dynamic learning rates $\eta_{t}=1/t^{1/2+\varepsilon}$ where $\varepsilon=0.1$. “Baseline” and “Clean” denote the standard online conformal prediction with noisy and clean labels, respectively. We employ LAC scores to generate prediction sets with error rate $\alpha=0.1$, using ResNet18 on ImageNet.
5 Experiments
Datasets and setup.
We evaluate the performance of NR-OCP under uniform label noise. The experiments include both constant $\eta=0.05$ and dynamic learning rates $\eta_{t}=1/t^{1/2+\varepsilon}$ with $\varepsilon=0.1$, following prior work (Angelopoulos et al., 2024)). We use CIFAR-100 (Krizhevsky, 2009) and ImageNet (Deng et al., 2009) dataset with synthetic label noise. On ImageNet, we use four pre-trained classifiers from TorchVision (Paszke et al., 2019) - ResNet18, ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and VGG16 (Simonyan & Zisserman, 2015). On CIFAR-100, we train these models for 200 epochs using SGD with a momentum of 0.9, a weight decay of 0.0005, and a batch size of 128. We set the initial learning rate as 0.1, and reduce it by a factor of 5 at 60, 120 and 160 epochs.
Conformal prediction algorithms. We apply LAC (Sadinle et al., 2019), APS (Romano et al., 2020b), RAPS (Angelopoulos et al., 2021) and SAPS (Huang et al., 2023) to generate prediction sets. A detailed description of these non-conformity scores is provided in Appendix B.4. For the evaluation metrics, we employ the long-run coverage gap (CovGap) and long-run size (Size). Formally, given a test dataset $\mathcal{I}_{test}$, these two metrics are defined as:
$\mathrm{CovGap}=\left|\frac{1}{|\mathcal{I}_{test}|}\sum_{t\in\mathcal{I}_{test}}\mathds{1}\left\{Y_{t}\in\mathcal{C}_{t}(X_{t})\right\}-(1-\alpha)\right|,$
$\mathrm{Size}=\frac{1}{|\mathcal{I}_{test}|}\sum_{t\in\mathcal{I}_{test}}|\mathcal{C}_{t}(X_{t})|.$
The long-run size measures the efficiency of prediction sets. Small prediction sets are preferred as they can provide precise predictions, thereby enabling accurate human decision-making in real-world scenarios (Cresswell et al., 2024).
Table 2: Performance of different methods using different models under uniform label noise with noise rate $\epsilon=0.05$ on CIFAR-100. “Baseline” denotes the standard online conformal prediction method. We employ LAC score to generate prediction sets with error rate $\alpha=0.1$ and use a constant learning rate $\eta=0.05$. $\downarrow$” indicates smaller values are better and Bold numbers are superior results.
Models CovGap(%) $\downarrow$ Size $\downarrow$
$\alpha=0.1$ $\alpha=0.05$ $\alpha=0.1$ $\alpha=0.05$
Baseline / Ours
ResNet18 2.744 / 0.289 1.900 / 0.122 31.61 / 22.78 56.84 / 47.31
ResNet50 2.700 / 0.378 1.389 / 0.289 33.23 / 19.21 59.60 / 50.24
DenseNet121 3.178 / 0.156 1.933 / 0.211 25.88 / 15.20 52.93 / 43.70
VGG16 1.833 / 0.411 1.011 / 0.067 52.33 / 42.11 72.69 / 68.03
Main results. In Table 1, we evaluate NR-OCP (Ours) against the standard online conformal prediction (Baseline) that updates the threshold with noisy labels (see Eq. (2) and Eq. (5)), employing a ResNet18 model. Due to space constraints, we report the average performance across four non-conformity score functions. The performance on each score function is provided in Appendix B.6. A salient observation is that NR-OCP eliminates the long-run coverage gap present in the baseline, achieving precise coverage guarantees while significantly improving the long-run efficiency of prediction sets. For example, on ImageNet with error rate $\alpha=0.1$, noise rate $\epsilon=0.15$ and dynamic learning rate, the baseline deviates from the target coverage level of $0.9$, exhibiting a coverage gap of 8.372% and an average set size of 171.2. In contrast, NR-OCP achieves a negligible coverage gap of 0.183% and a prediction set size of 13.10.
Moreover, following prior work (Angelopoulos et al., 2024), we demonstrate the dynamics of coverage rate in Figure 2. We use the LAC score and a ResNet18 model on ImageNet, with error rate $\alpha=0.1$ and noise rate $\epsilon=0.05$. The results show that NR-OCP successfully converges to the desired coverage rate of $1-\alpha$ on both constant and dynamic learning rates. Overall, NR-OCP demonstrates superior performance over the baseline by achieving both precise coverage guarantee and improved prediction set efficiency.
NR-OCP is effective on different model architectures. In Table 2, we show that NR-OCP is effective on a diverse range of model architectures. The experiments are conducted on CIFAR-100, using LAC score to generate prediction sets with error rate $\alpha=0.1$ and noise rate $\epsilon=0.05$, under constant learning rate schedule. In particular, our method consistently outperforms the baseline across various models. For example, on DenseNet121 with error rate $\alpha=0.1$, the baseline exhibits a coverage gap of $3.178\%$ and size of $25.88$, while NR-OCP achieves a smaller coverage gap of $0.156\%$ and a reduced size of $15.20$. Additional results on these models is available in Appendix B.7.
@@ 6 Conclusion
In this work, we investigate the robustness of online conformal prediction under uniform label noise with a known noise rate, in both constant and dynamic learning rate schedule. Our theoretical analysis shows that the presence of label noise causes a deviation between the actual and desired mis-coverage rate $\alpha$, with higher noise rates resulting in larger gaps. This gap can lead to either overestimated or underestimated coverage guarantees, which depend on the size of the prediction sets. To address this issue, we propose Noise Robust Online Conformal Prediction (dubbed NR-OCP) which updates the threshold using the proposed robust pinball loss. We establish that this loss is equivalent to the pinball loss under clean labels in expectation. We prove that NR-OCP eliminates the coverage gap caused by the label noise, in both constant and dynamic learning rate schedules, achieving a convergence rate of $\mathcal{O}(T^{-1/2})$ for both empirical and expected coverage errors under the uniform label noise. Extensive experiments verify that our method achieves superior performance over the baseline by achieving both precise coverage rates and improved set efficiency across various settings, including various model architectures, error rates, noise rates, and non-conformity scores. We hope the theoretical framework in this work can inspire more specifically designed methods for online conformal prediction under label noise.
Limitation. In this work, our analysis and method are limited to the setting of uniform noisy labels with a known noise rate. We believe it will be interesting to develop online conformal prediction algorithms that are robust to various types of label noise, in the future.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Related work
Conformal prediction (Papadopoulos et al., 2002; Vovk et al., 2005) is a statistical framework for uncertainty qualification. In the literature, many conformal prediction methods have been proposed across various domains, such as regression (Lei & Wasserman, 2014; Romano et al., 2020a), image classification (Angelopoulos et al., 2021; Huang et al., 2024; Liu et al., 2024), outlier detection (Guan & Tibshirani, 2022; Bates et al., 2023; Liang et al., 2024), and large language models (Gui et al., 2024; Cherian et al., 2024). Conformal prediction is also deployed in other real-world applications, such as human-in-the-loop decision-making (Cresswell et al., 2024), automated vehicles (Bang et al., 2024), and scientific computation (Moya et al., 2024). In what follows, we introduce the most related works in two settings: online learning and noise robustness.
Online conformal prediction. Conventional conformal prediction algorithms provide coverage guarantees under the assumption of data exchangeability. However, in real-world online scenarios, the data distribution may evolve over time, violating the exchangeability assumption (Zhang et al., 2018; Zhao et al., 2022). To address this challenge, recent research develops online conformal prediction methods that adaptively construct prediction sets to accommodate distribution shift (Gibbs & Candes, 2021; Feldman et al., 2023; Bhatnagar et al., 2023; Angelopoulos et al., 2024; Gibbs & Candès, 2024; Hajihashemi & Shen, 2024). Building on online convex optimization techniques (Anderson, 2008; Moore et al., 2011; Hazan et al., 2016; Singh & Thurman, 2019; Hoi et al., 2021; Orabona, 2019), these methods employ online gradient descent with pinball loss to provably achieve desired coverage under arbitrary distributional changes (Gibbs & Candes, 2021). Still, these algorithms typically assume perfect label accuracy, an assumption that rarely holds in practice, given the prevalence of noisy labels in online learning (Ben-David et al., 2009; Natarajan et al., 2013; Wu et al., 2024). In this work, we theoretically show that label noise can significantly affect the long-run mis-coverage rate through biased pinball loss gradients, leading to either inflated or deflated coverage guarantee.
Noise-robust conformal prediction. The issue of label noise has been a common challenge in machine learning with extensive studies (Xia et al., 2019; Wei et al., 2020; Chen et al., 2021; Wei et al., 2021; Wu et al., 2021; Li et al., 2021; Zhu et al., 2022; Wei et al., 2023; Chen et al., 2024; Gao et al., 2024). In the context of conformal prediction, recent works develop noise-robust conformal prediction algorithms for both uniform noise (Penso & Goldberger, 2024) and noise transition matrix (Sesia et al., 2024). The most relevant work (Einbinder et al., 2024) shows that online conformal prediction maintains valid coverage when noisy scores stochastically dominate clean scores. Our analysis extends this work by removing the assumption on noisy and clean scores, offering a more general theoretical framework for understanding the impact of uniform label noise.
@@ B.1 Regret analysis
As Bhatnagar et al. (2023) demonstrates, regret serves as a helpful performance measure alongside coverage. In particular, it can identify algorithms that achieve valid coverage guarantees through impractical means. For example, prediction sets that alternate between empty and full sets with frequencies $\{\alpha,1-\alpha\}$ satisfy coverage bounds on any distribution but have linear regret on simple distributions (see detailed proof in Appendix A.2 of Bhatnagar et al. (2023)). Drawing from standard online learning theory (see Theorem 2.13. of Orabona (2019)), we analyze the regret bound for our method. Let $\tau^{*}:=\arg\min_{\hat{\tau}}\sum_{t=1}^{T}\tilde{l}_{1-\alpha}(\tau^{*},\tilde{S}_{t},\{S_{t,y}\}_{y=1}^{K})$, and define the regret as
@@
This leads to the following regret bound (the proof is provided in Appendix E):
@@
This linear regret aligns with the observation in Angelopoulos et al. (2024): the constant learning rate schedule, while providing valid coverage on average, lead to significant temporal variability in Coverage($\hat{\tau}_{t}$) (see Proposition 1 in Angelopoulos et al. (2024)). The linear regret bound provides additional theoretical justification for this instability.
@@ B.2 Convergence analysis of $\hat{\tau}_{t}$
We analyze the convergence of our method toward global minima by combining the convergence analysis of SGD (see Theorem 5.3. in Garrigos & Gower (2023)) and self-calibration inequality of pinball loss (see Theorem 2.7. in Steinwart & Christmann (2011)). Following Steinwart & Christmann (2011), we first establish assumptions on the distribution of $S$. Let $R:=2S-1\in[0,1]$, with $\gamma$ denoting the $1-\alpha$ quantile of $R$. This indicates that $\gamma=2\tau-1$, where $\tau$ is the $1-\alpha$ quantile of $S$. We make the following assumption:
@@
Furthermore, let $\beta=b/(q-1)$ and $\delta=\beta(2\varepsilon_{0})^{q-1}$. This leads to the following proposition (the proof is provided in Appendix F):
@@ B.3 Why does standard online conformal prediction rarely exhibit under-coverage under label noise?
As established in Proposition 3.1 and 4.1, the label noise introduces a gap of $\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right)$ between the actual mis-coverage rate and desired mis-coverage rate $\alpha$. This result indicates that when the prediction sets are sufficiently large such that $(1-\alpha)\leq\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]$, a high noise rate $\epsilon$ decreases this upper bound, resulting in under-coverage prediction sets, i.e., $\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}\geq\alpha$. However, this scenario only arises when $\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\geq K(1-\alpha)$. To illustrate, considering CIFAR-100 with $K=100$ classes and error rate $\alpha=0.1$, under-coverage would require $\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\geq 90$, a condition that remains improbable even with random predictions. Moreover, in such extreme cases, as $(1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]$ approaches 0, the coverage gap becomes negligible, resulting in coverage rates that approximate the desired $1-\alpha$. We verify this empirically in Figure 3, where we implement online conformal prediction using an untrained, randomly initialized ResNet18 model on CIFAR-10 and CIFAR-100 with noise rates $\epsilon=0.05,0.1,0.15$. The results confirm that even in this extreme scenario, the coverage gap remains negligible, with coverage rates approaching the desired 0.9.

### v3

@@ 5.2 Main results
Table 1: Average performance of different methods under uniform noisy labels across 4 score functions, using ResNet18. The performance on each score function is provided in Appendix B.5. “Baseline” denotes the ACI with standard pinball loss. We include two learning rate schedules: constant learning rate $\eta=0.05$ and dynamic learning rates $\eta_{t}=1/t^{1/2+\varepsilon}$ where $\varepsilon=0.1$. $\downarrow$” indicates smaller values are better and Bold numbers are superior results.
@@
Constant
$\epsilon=0.05$ Baseline 3.942 2.867 11.93 30.66 4.163 3.289 101.0 231.7
@@
Dynamic
$\epsilon=0.05$ Baseline 4.106 2.780 6.527 21.97 4.498 2.584 31.15 128.1
@@ 6 Conclusion
In this work, we investigate the robustness of online conformal prediction under uniform label noise with a known noise rate, in both constant and dynamic learning rate schedules. Our theoretical analysis shows that the presence of label noise causes a deviation between the actual and desired mis-coverage rate $\alpha$, with higher noise rates resulting in larger gaps. To address this issue, we propose a novel loss function robust pinball loss, which provides an unbiased estimate of clean pinball loss without requiring ground-truth labels. We theoretically establish that this loss is equivalent to the pinball loss under clean labels in expectation. In our theoretical analysis, we show that robust pinball loss enables online conformal prediction to eliminate the coverage gap caused by the label noise, achieving a convergence rate of $\mathcal{O}(T^{-1/2})$ for both empirical and expected coverage errors under uniform label noise. This loss offers a general solution to the uniform label noise, and is complementary to existing online conformal prediction methods. Extensive experiments demonstrate that the proposed loss enhances the noise robustness of various online conformal prediction methods by eliminating the coverage gap caused by the label noise. Notably, our loss function is effective across different label noise, error rate, non-conformity score functions, and model architectures.
Limitation.
As the first step to explore the label noise issue in online conformal prediction, our analysis and method are limited to the setting of uniform noisy labels with a known noise rate. We believe it will be interesting to develop online conformal prediction algorithms that are robust to various types of label noise with fewer assumptions in the future.
@@ Appendix A Related work
Conformal prediction [38, 5] is a statistical framework for uncertainty qualification. In the literature, many conformal prediction methods have been proposed across various domains, such as regression [39, 40], image classification [35, 41, 42], outlier detection [43, 44, 45], and large language models [46, 47]. Conformal prediction is also deployed in other real-world applications, such as human-in-the-loop decision-making [37], automated vehicles [48], and scientific computation [49]. In what follows, we introduce the most related works in two settings: online learning and noise robustness.
Online conformal prediction. Conventional conformal prediction algorithms provide coverage guarantees under the assumption of data exchangeability. However, in real-world online scenarios, the data distribution may evolve over time, violating the exchangeability assumption [50, 51]. To address this challenge, recent research develops online conformal prediction methods that adaptively construct prediction sets to accommodate distribution shift [9, 10, 11, 12, 13, 14]. Building on online convex optimization techniques [52, 53, 54, 55, 56, 28], these methods employ online gradient descent with pinball loss to provably achieve desired coverage under arbitrary distributional changes [9]. Still, these algorithms typically assume perfect label accuracy, an assumption that rarely holds in practice, given the prevalence of noisy labels in online learning [15, 16, 17]. In this work, we theoretically show that label noise can significantly affect the long-run mis-coverage rate through biased pinball loss gradients, leading to either inflated or deflated coverage guarantee.
Noise-robust conformal prediction. The issue of label noise has been a common challenge in machine learning with extensive studies [57, 27, 58, 59, 60, 61, 62, 63, 64, 65]. In the context of conformal prediction, recent works develop noise-robust conformal prediction algorithms for both uniform noise [23] and noise transition matrix [24]. The most relevant work [18] shows that online conformal prediction maintains valid coverage when noisy scores stochastically dominate clean scores. Our analysis extends this work by removing the assumption on noisy and clean scores, offering a more general theoretical framework for understanding the impact of uniform label noise.
@@ B.1 Regret analysis
As [11] demonstrates, regret serves as a helpful performance measure alongside coverage. In particular, it can identify algorithms that achieve valid coverage guarantees through impractical means. For example, prediction sets that alternate between empty and full sets with frequencies $\{\alpha,1-\alpha\}$ satisfy coverage bounds on any distribution but have linear regret on simple distributions (see detailed proof in Appendix A.2 of [11]). Drawing from standard online learning theory (see Theorem 2.13. of [28]), we analyze the regret bound for our method. Let $\tau^{*}:=\arg\min_{\hat{\tau}}\sum_{t=1}^{T}\tilde{l}_{1-\alpha}(\tau^{*},\tilde{S}_{t},\{S_{t,y}\}_{y=1}^{K})$, and define the regret as
@@
This leads to the following regret bound (the proof is provided in Appendix LABEL:appendix:proof:prop:regret_analysis):
@@
This linear regret aligns with the observation in [12]: the constant learning rate schedule, while providing valid coverage on average, lead to significant temporal variability in Coverage($\hat{\tau}_{t}$) (see Proposition 1 in [12]). The linear regret bound provides additional theoretical justification for this instability.
@@ B.2 Convergence analysis of $\hat{\tau}_{t}$
We analyze the convergence of our method toward global minima by combining the convergence analysis of SGD (see Theorem 5.3 in [29]) and self-calibration inequality of pinball loss (see Theorem 2.7. in [66]). Following [66], we first establish assumptions on the distribution of $S$. Let $R:=2S-1\in[0,1]$, with $\gamma$ denoting the $1-\alpha$ quantile of $R$. This indicates that $\gamma=2\tau-1$, where $\tau$ is the $1-\alpha$ quantile of $S$. We make the following assumption:
@@
Furthermore, let $\beta=b/(q-1)$ and $\delta=\beta(2\varepsilon_{0})^{q-1}$. This leads to the following proposition (the proof is provided in Appendix LABEL:proof:prop:convergence_global_minima):
@@ B.3 Why does standard online conformal prediction rarely exhibit under-coverage under label noise?
As established in Proposition 3.1, the label noise introduces a gap of $\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right)$ between the actual mis-coverage rate and desired mis-coverage rate $\alpha$. This result indicates that when the prediction sets are sufficiently large such that $(1-\alpha)\leq\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]$, a high noise rate $\epsilon$ decreases this upper bound, resulting in under-coverage prediction sets, i.e., $\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}\geq\alpha$. However, this scenario only arises when $\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\geq K(1-\alpha)$. To illustrate, considering CIFAR-100 with $K=100$ classes and error rate $\alpha=0.1$, under-coverage would require $\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\geq 90$, a condition that remains improbable even with random predictions. Moreover, in such extreme cases, as $(1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]$ approaches 0, the coverage gap becomes negligible, resulting in coverage rates that approximate the desired $1-\alpha$. We verify this empirically in Figure 3, where we implement online conformal prediction using an untrained, randomly initialized ResNet18 model on CIFAR-10 and CIFAR-100 with noise rates $\epsilon=0.05,0.1,0.15$. The results confirm that even in this extreme scenario, the coverage gap remains negligible, with coverage rates approaching the desired 0.9.